OpenStudy (anonymous):

Let a and b be the roots of x^2 - 6x - 2 =0 with a>b . if cn = a^n - b^n for n>=1 then the value of (c10-2c8)/2c9

OpenStudy (anonymous):

@dumbcow @satellite73 @experimentX

OpenStudy (experimentx):

??

OpenStudy (anonymous):

if a b be the roots then a*b= 2 and a+b=6

OpenStudy (anonymous):

ab = -2

OpenStudy (anonymous):

ooops typo

OpenStudy (experimentx):

(c10-2c8)/2c9 ??

OpenStudy (anonymous):

if cn = a^n - b^n

OpenStudy (anonymous):

now try to love \[a ^{10}-b ^{10}= (a ^{5}-b ^{5}) (a ^{5}+b ^{5})\] like this

OpenStudy (anonymous):

hey..hw to get rid off ^5

OpenStudy (anonymous):

you will see when you will take \[a ^{4}+b ^{4}\] common out of numerator just solve it like i said and you will see

OpenStudy (experimentx):

try playing with this.